# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

### v4

@@ 3.2.3 Optimization Design
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
@@
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Table.2.
The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.5, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
@@ 3.2.4 Physical Integrations and Experiments
Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation. Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
@@ 3.2.5 Count Rate and Timing Jitter
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
@@ 4 Conclusion
In summary, the enhanced single-photon detection (ESPD) framework is established to reformulate photon detection as an iteratively enhanced quantum-information-processing task composed of state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a fixed device-level operation. Within this system-level theoretical framework, detection performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
Numerical simulations, grounded in experimentally reported parameters and commercially available photonic components, demonstrate that the ESPD framework can substantially enhance the effective detection efficiency (DE) (from 59% to 93%) while strongly suppressing the effective dark count rate (DCR) (from $10^{-2}$ to $10^{-9}$) even when initialized with a conventional room-temperature SPD reported two decades ago. The resulting effective performance metrics are comparable to those of state-of-the-art superconducting single-photon detectors, significantly relaxing the minimal tolerable channel transmission rate in quantum communication applications.
Beyond performance enhancement, the ESPD framework makes explicit the fundamental trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure requirements with system-level integration complexity. Although the realization of large-scale ESPD architectures would require further advances in device integration and experimental implementation, the theoretical formulation and parameter studies presented here establish a concrete foundation for the framework and provide concrete guidance for future experimental efforts.
In essence, this work establishes a conceptually complete and theoretically grounded framework for high-performance single-photon detection at room temperature. It demonstrates that intrinsic device-level limitations need not be overcome solely through component replacement, but can instead be effectively compensated by system-level architectural design. As such, the ESPD framework offers a viable pathway toward next-generation photon detection and provides a general methodology for rethinking performance limits in quantum technologies under practical constraints.
Acknowledgements
The author acknowledges the use of Gemini and ChatGPT (accessed December 2025) to, and only to, support language polishing, flow paragraphs, and correct grammar, to improve readability. All AI-generated suggestions were reviewed, revised, and approved by the authors, who take full responsibility for the accuracy and integrity of the work. All insight, results, methods, analysis, and experiments are ordinarily created and owned by the author.

### v5

@@ 3.3.3 Optimization Design
The inherent exponential resource scaling discussed in Section 3.3.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Minimizing the iteration depth ($L$) provides significant savings in component count and system complexity due to the exponent dependence. As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.10, with a suitably selected $n,k$ on each level, the convergence can be accelerated, and the intermediate-level state can be optimized, which is crucial for minimizing resource consumption.
Table 7: Performance evolution of the ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n_{s},k_{s})$. Settings are the same as in Section 2.4
@@
(a) DE $\eta_{s}$ vs. level $s$.
(b) DCR $d_{s}$ vs. level $s$.
Figure 10: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different values of $n,k$. Para 3 is the same as in Table.2.
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^5] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical complexity: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
@@ 3.3.4 Physical Integrations and Experiments
Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation. Reassuringly, as established in Section 3.2, the scheme relies only on readily available quantum optical components, suggesting no fundamental technological barriers prevent its realization. However, the realization of the large-scale iterative circuit demands significant engineering efforts, particularly concerning the precise integration and synchronization of components. Furthermore, the seamless integration of the ESPD into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
@@ 3.3.5 Count Rate and Timing Jitter
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and count rate are primarily determined by the intrinsic temporal response of the baseline SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, such that the architectural integration introduces negligible temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics as well.
@@ 4 Conclusion
In summary, we have established the enhanced single-photon detection (ESPD) framework, which reconceptualizes photon detection as an iteratively enhanced quantum-information-processing task that integrates state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a static device-level operation. Within this system-level theoretical framework, the detection performance is formulated as a non-linear dynamical system, where the effective metrics are progressively optimized through structured quantum processing using exclusively non-superconducting, room-temperature components.
Analytical approximations, Monte Carlo analysis, and numerical simulations, grounded in experimentally reported parameters and commercial photonic specifications, demonstrate that the ESPD framework drives detection performance toward the high-performance basin of attraction, provided the initial conditions lie within a suitable operating range. Consequently, the performance of traditional single-photon detectors (SPDs) can be substantially enhanced: the effective detection efficiency (DE) can be boosted (e.g., from 59% to 93%) while the effective dark count rate (DCR) can be suppressed by several orders of magnitude (e.g., from $10^{-2}$ to $10^{-9}$), to the level comparable to state-of-the-art superconducting SPDs, resulting in the significantly expansion of the operational boundaries of quantum communication by relaxing the minimal tolerable channel transmission rate.
Beyond performance enhancement, the ESPD framework establishes a trade-off between resource overhead and environmental constraints, effectively translating the demand for cryogenic cooling during operation into a one-time integration complexity. While large-scale implementation poses engineering challenges in component synchronization and integration density, the presenting theoretical formulations and technological feasibility studies establish a rigorous foundation and provide a roadmap for future experimental realizations.
In essence, this work provides a theoretically grounded framework for high-performance photon detection at room temperature. It demonstrates that intrinsic device-level limitations can be circumvented through system-level architectural design rather than reliance exclusively on material breakthroughs. By establishing the dynamical evolution of the ESPD paradigm, this work not only offers a viable pathway toward next-generation room-temperature photonic devices but also introduces a general methodology for rethinking performance limits in quantum technologies under real-world constraints, potentially influencing a wide range of quantum applications.